# Constraining primordial black hole abundance with Insight-HXMT

### v2

@@ 3 Constrains with Insight-HXMT data
The research in ref. [27] indicates that when considering the impact of positron annihilation on the abundance of PBHs, both the extended mass function and the spin of the PBHs lead to stronger constraints. When the mass distribution function is assumed to be log-normal, $f_{\mathrm{PBH}}$ changes little as $\tau$ increases before exceeding a certain threshold; after the threshold is surpassed, $f_{\mathrm{PBH}}$ starts to decay rapidly. When considering black hole spin, $f_{\mathrm{PBH}}$ exhibits a trend of continuously accelerating decay as the spin parameter $\alpha$ increases. Therefore, in our work, to obtain more conservative and broadly applicable constraint results, we consider only the monochromatic, non-spinning PBHs, which represent the most general case.
3.1 Constraints derived directly from observational data
@@
Figure 2: The constraint results in our work under three different scenarios: directly using observational data (blue solid line), incorporating the astrophysical background model (blue dashed line), and applying the power-law model fitted to the observational data (blue dotted-dashed line). These results are compared with existing constraints: magenta from [30], red from [31], cyan from [28], brown from [27] and green from [32].
As shown in figure 2, when we directly use observational data to constrain the abundance of PBHs, our results are significantly weaker compared to previous studies [30, 31, 19, 27, 32]. This discrepancy primarily arises from the limited availability of high-energy observational data in our analysis, as well as the absence of background subtraction for diffuse sources. These background sources contribute significantly to the measured flux, leading to a considerable underestimation of our constraints.
In figure 2, our result exhibits a simple linear relationship, which contrasts with the irregular shapes of the other curves. This is primarily due to two factors. First, with the exception of Iguaz et al. [27], which adopts a similar approach to ours, all other studies in figure 2 consider only the contribution from the Galactic photon flux component, ${\mathrm{d}\phi_{\gamma}^{\text{gal}}}/{\mathrm{d}E}$, in Eq. (2.10). Second, our analysis considers only the theoretical upper energy limit of Insight-HXMT, approximately 250 keV, while the data of ref. [27] are derived from multiple instruments with upper energy limits exceeding 10 MeV.
From figure 2, it is evident that within the energy range of Insight-HXMT, the contributions from ${\mathrm{d}\phi_{\text{ann}}^{\text{ext}}}/{\mathrm{d}E}$ and ${\mathrm{d}\phi_{\text{ann}}^{\text{gal}}}/{\mathrm{d}E}$ in Eq. (2.10) dominate the photon flux. In this regime, as $M_{\mathrm{PBH}}$ varies, the spectral shapes of ${\mathrm{d}\phi_{\text{ann}}^{\text{ext}}}/{\mathrm{d}E}$ and ${\mathrm{d}\phi_{\text{ann}}^{\text{gal}}}/{\mathrm{d}E}$ change mainly in terms of slope, resulting in minimal shifts in the optimal data points used to constrain $f_{\mathrm{PBH}}$, which consistently lies near the detector’s upper energy boundary. This leads to a nearly linear relationship between $M_{\mathrm{PBH}}$ and $f_{\mathrm{PBH}}$ on a logarithmic scale.
@@
3.2 Situation with astrophysical background
To mitigate the influence of diffuse background sources and thus obtain more accurate results, we adopted the ABM proposed in ref. [33]. Since the dominant contributors to the diffuse X-ray background are blazars and AGNs, the model combines AGN and blazar emissions using a double power-law spectrum to fit the observational data. Its specific form is:
${\psi^{\text{ABM}}}=\frac{\mathrm{d}\phi^{\text{ABM}}}{\mathrm{d}E}=\frac{A}{\left(E/E_{b}\right)^{n_{1}}+\left(E/E_{b}\right)^{n_{2}}},$ (3.1)
where the best-fit parameters are: $E_{b}=35.6966~{}\text{keV}$, $A=0.0642~{}\text{keV}^{-1}~{}\text{s}^{-1}~{}\text{cm}^{-2}~{}\text{sr}^{-1}$, $n_{1}=1.4199$, and $n_{2}=2.8956$. The observational data used to derive this fitting formula primarily originate from Solar Maximum Mission (SMM), Nagoya balloon, and High Energy Astrophysical Observatory (HEAO).
@@ 3.2 Situation with astrophysical background
$\chi^{2}=\sum\left[\frac{\psi^{\text{data }}-\psi^{\mathrm{PBH}}-\psi^{\mathrm{ABM}}}{\sigma^{\text{data }}}\right]^{2}.$ (3.2)
@@
Here $\psi^{\text{data}}$ represents the differential photon flux per solid angle from diffuse X-ray background data, $\sigma^{\text{data}}$ is the associated uncertainty in $\psi^{\text{data}}$. We sum over all energy ranges of the data to calculate the ${\chi}^{2}$ value. To determine the 95% confidence level bound on $f_{\mathrm{PBH}}$, we calculate the ${\chi}^{2}$ value for each PBH mass M, requiring $\chi^{2}-\chi_{\min}^{2}\leq 3.84$. The constraints considering the astrophysical background model are shown in figure 2.
@@
It is clear that when we remove the influence of the astrophysical background, the constraints on the PBH abundance are significantly improved. In the mass range of $2\times 10^{16}$$5\times 10^{17}$ g. The constraints are improved by approximately 1.5 orders of magnitude compared to the results obtained by directly using the observational data. While our results have been significantly improved, they only surpass some previous studies in the mass range above $10^{17}$ g. We attribute the weaker constraints to the lack of high-energy observational data. Furthermore, the method of using a double power-law spectrum to fit the observational data, which combines AGN and blazar emissions, to obtain the astrophysical background model, still appears somewhat rough and not sufficiently accurate. In future work, we plan to adopt the template analysis method from ref. [32] to analyze the Insight-HXMT observational data, in order to accurately determine the contribution of each background source to the diffuse X-ray background. This will further improve our constraint results.
3.3 Extend the energy range of the diffuse X-ray background
@@
The observational data we used from ref. [12] spans an energy range of 1.5–120 keV, although the full observational range of Insight-HXMT extends from 1 to 250 keV. If data above 120 keV become available, Insight-HXMT could potentially provide more stringent constraint results. Considering this possibility, we performed a preliminary estimate of the constraints over the full 1–250 keV range.
@@
Ref. [12] used a power-law spectrum with an exponential cutoff to fit the observational data. Based on the best-fit parameters provided in the reference, we can model the diffuse X-ray background as follows:
$\frac{\mathrm{d}\phi_{\gamma}}{\mathrm{d}E}=9.57~{}E^{-1.402}~{}e^{\frac{E}{55\mathrm{~{}keV}}}\mathrm{~{}cm}^{-2}\mathrm{~{}s}^{-1}\mathrm{~{}keV}^{-1}\mathrm{~{}sr}^{-1}.$ (3.3)
@@
We assume that this power-law (PL) model remains valid up to the upper observational limit of Insight-HXMT’s energy range. Therefore, within the 1–250 keV range, we require that the total photon flux from PBHs not exceed the photon flux of the diffuse X-ray background predicted by this model. The final constraint results also are presented in figure 2.
@@
The constraint results, obtained by extrapolating the diffuse X-ray background model fitted to the Insight-HXMT observational data to the upper energy limit of the detector, are approximately 0.5 orders of magnitude stronger than those based on the astrophysical background model. In this case, our observational results surpass those of most previous studies, providing stronger constraints than all others in the mass range above $10^{17}$ g. The exclusion limit for PBHs as dark matter has also been extended to $4\times 10^{17}~{}\mathrm{g}$, which aligns with the best upper limit presented in ref. [32]. Due to the lack of sufficient observational data, we are unable to employ ${\chi}^{2}$ estimation method to account for the influence of background diffuse sources. Therefore, we believe that with adequate observational data, Insight-HXMT will be capable of providing even more stringent constraint results.
4 Conclusion
@@
In this work, we primarily utilize observational data from Insight-HXMT to improve the existing constraints on the abundance of PBHs in the mass range of $2\times 10^{16}$$5\times 10^{17}$ g. We first consider the photon flux emitted by PBHs within this mass range, incorporating contributions from both Galactic and extragalactic sources. In this range, the photon flux generated by positron annihilation due to Hawking radiation cannot be neglected, particularly when the black hole temperature falls below 511 keV, where annihilation photons dominate. Therefore, we account for both the photon flux directly produced by Hawking radiation and the photons resulting from positron annihilation.
@@
First, when we directly use the observational data from Insight-HXMT to derive the constraint results, due to the lack of high-energy observational data and the absence of background diffuse source subtraction, our results are weaker than those of previous studies. When we introduce the astrophysical background model, the constraints improve by approximately 1.5 orders of magnitude relative to the first scenario. Although the constraint results have been improved significantly, we still do not have an advantage compared to previous studies.
@@
In our analysis, the energy range of the data used spans from 0 to 120 keV, whereas the actual upper energy limit of Insight-HXMT observations extends up to 250 keV. Although we used the PL model to estimate the constraints based on data from 0 to 250 keV, the results remain less precise. Moreover, due to the lack of actual observational data, we are unable to remove the impact of diffuse background sources. This limitation prevents us from obtaining tighter constraints when using the power-law model. We need more observational data, especially in the 120 keV to 250 keV energy range, to improve our results. Furthermore, we applied an existing astrophysical background model without conducting a detailed analysis or disentangling individual sources within the data. Consequently, the results obtained are not the most accurate. We anticipate that by addressing these issues in future work, we will be able to achieve more accurate and robust results.
[…]
Acknowledgments
We thank Sai Wang, Yichao Li, Chen Zhang, and Rui Huang for helpful discussions. Special thanks are given to Rui Huang for providing the observational data. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grant Nos. 12473001 and 11975072), and the National 111 Project (Grant No. B16009).

### v3

@@ III Constraints with Insight-HXMT data
The research in Ref. Iguaz:2021irx indicates that when considering the impact of positron annihilation on the abundance of PBHs, both the extended mass function and the spin of the PBHs lead to stronger constraints. When the mass distribution function is assumed to be log-normal, $f_{\mathrm{PBH}}$ exhibits negligible variation with increasing width parameter $\sigma$ until $\sigma$ exceeds a critical threshold; after exceeding the threshold, $f_{\mathrm{PBH}}$ starts to decay rapidly. When considering black hole spin, $f_{\mathrm{PBH}}$ exhibits a trend of continuously accelerating decay as the dimensionless spin parameter $\alpha$ increases. Therefore, in our work, to obtain more conservative and broadly applicable constraint results, we consider only the monochromatic, non-spinning PBHs, which represent the most general case.
III.1 Constraints derived directly from observational data
@@
Figure 2: The constraint results in our work under three different scenarios: directly using observational data (blue solid line), incorporating the astrophysical background model (blue dashed line), and applying the power-law model fitted to the observational data (blue dotted-dashed line). These results are compared with existing constraints: magenta from Laha:2019ssq, red from Laha:2020ivk, cyan from Coogan:2020tuf, brown from Iguaz:2021irx and green from Berteaud:2022tws.
As shown in Fig. 2, when we directly use observational data to constrain the abundance of PBHs, our results are significantly weaker compared to previous studies Laha:2019ssq; Laha:2020ivk; Coogan:2021sjs; Iguaz:2021irx; Berteaud:2022tws. This discrepancy primarily arises from the limited availability of high-energy observational data in our analysis, as well as the absence of background subtraction for diffuse sources. These background sources contribute significantly to the measured flux, leading to a considerable underestimation of our constraints.
In Fig. 2, our result exhibits a simple linear relationship, which contrasts with the irregular shapes of the other curves. This is primarily due to two factors. First, with the exception of Iguaz et al. (2021), which adopts a similar approach to ours, all other studies in Fig. 2 consider only the contribution from the Galactic photon flux component, ${\mathrm{d}\phi_{\gamma}^{\text{gal}}}/{\mathrm{d}E}$, in Eq. (10). Second, our analysis considers only the theoretical upper energy limit of Insight-HXMT, approximately 250 keV, while Iguaz’s data are derived from multiple instruments with upper energy limits exceeding 10 MeV.
From Fig. 2, it is evident that within the energy range of Insight-HXMT, the contributions from ${\mathrm{d}\phi_{\text{ann}}^{\text{ext}}}/{\mathrm{d}E}$ and ${\mathrm{d}\phi_{\text{ann}}^{\text{gal}}}/{\mathrm{d}E}$ in Eq. (10) dominate the photon flux. In this regime, as $M_{\mathrm{PBH}}$ varies, the spectral shapes of ${\mathrm{d}\phi_{\text{ann}}^{\text{ext}}}/{\mathrm{d}E}$ and ${\mathrm{d}\phi_{\text{ann}}^{\text{gal}}}/{\mathrm{d}E}$ change mainly in terms of slope, resulting in minimal shifts in the optimal data point used to constrain $f_{\mathrm{PBH}}$, which consistently lies near the detector’s upper energy boundary. This leads to a nearly linear relationship between $M_{\mathrm{PBH}}$ and $f_{\mathrm{PBH}}$ on a logarithmic scale.
@@
III.2 Situation with astrophysical background
To mitigate the influence of diffuse background sources and thus obtain more accurate results, we adopted the ABM proposed in Ref. Ballesteros:2019exr. Since the dominant contributors to the diffuse X-ray background are blazars and AGNs, the model combines AGN and blazar emissions using a double power-law spectrum to fit the observational data. Its specific form is:
${\psi^{\text{ABM}}}=\frac{\mathrm{d}\phi^{\text{ABM}}}{\mathrm{d}E}=\frac{A}{\left(E/E_{b}\right)^{n_{1}}+\left(E/E_{b}\right)^{n_{2}}},$ (11)
where the best-fit parameters are: $E_{b}=35.6966~\text{keV}$, $A=0.0642~\text{keV}^{-1}~\text{s}^{-1}~\text{cm}^{-2}~\text{sr}^{-1}$, $n_{1}=1.4199$, and $n_{2}=2.8956$. The observational data used to derive this fitting formula primarily originate from Solar Maximum Mission (SMM), Nagoya balloon, and High Energy Astronomy Observatory (HEAO).
@@ III.2 Situation with astrophysical background
$\chi^{2}=\sum\left[\frac{\psi^{\text{data }}-\psi^{\mathrm{PBH}}-\psi^{\mathrm{ABM}}}{\sigma^{\text{data }}}\right]^{2}.$ (12)
@@
Here $\psi^{\text{data}}$ represents the differential photon flux per solid angle from diffuse X-ray background data, $\sigma^{\text{data}}$ is the associated uncertainty in $\psi^{\text{data}}$. We sum over all energy ranges of the data to calculate the ${\chi}^{2}$ value. To determine the 95% confidence level bound on $f_{\mathrm{PBH}}$, we calculate the ${\chi}^{2}$ value for each PBH mass M, requiring $\chi^{2}-\chi_{\min}^{2}\leq 3.84$. The constraints considering the astrophysical background model are shown in Fig. 2.
@@
It is clear that when we remove the influence of the astrophysical background, the constraints on the PBH abundance are significantly improved. In the mass range $2\times 10^{16}$$5\times 10^{17}$ g, the constraints are improved by approximately 1.5 orders of magnitude compared to the results obtained by directly using the observational data. While our results have been significantly improved, they only surpass some previous studies in the mass range above $10^{17}$ g. We attribute the weaker constraints to the lack of high-energy observational data. Furthermore, the method of using a double power-law spectrum to fit the observational data, which combines AGN and blazar emissions, to obtain the astrophysical background model, still appears somewhat rough and not sufficiently accurate. In future work, we plan to adopt the template analysis method from Ref. Berteaud:2022tws to analyze the Insight-HXMT observational data, in order to accurately determine the contribution of each background source to the diffuse X-ray background. This will further improve our constraint results.
III.3 Extending the energy range of the diffuse X-ray background
@@
The observational data we used from Ref. Huang:2022guk spans an energy range of 1.5–120 keV, although the full observational range of Insight-HXMT extends from 1 to 250 keV. If data above 120 keV become available, Insight-HXMT could potentially provide more stringent constraint results. Considering this possibility, we performed a preliminary estimate of the constraints over the full 1–250 keV range.
@@
Ref. Huang:2022guk used a power-law spectrum with an exponential cutoff to fit the observational data. Based on the best-fit parameters provided in the reference, we can model the diffuse X-ray background as follows:
$\frac{\mathrm{d}\phi_{\gamma}}{\mathrm{d}E}=9.57~E^{-1.402}~e^{-\frac{E}{55\mathrm{~keV}}}\mathrm{~cm}^{-2}\mathrm{~s}^{-1}\mathrm{~keV}^{-1}\mathrm{~sr}^{-1}.$ (13)
@@
We assume that this power-law (PL) model remains valid up to the upper observational limit of Insight-HXMT’s energy range. Therefore, within the 1–250 keV range, we require that the total photon flux from PBHs not exceed the photon flux of the diffuse X-ray background predicted by this model. The final constraint results also are presented in Fig. 2.
@@
The constraint results, obtained by extrapolating the diffuse X-ray background model fitted to the Insight-HXMT observational data to the upper energy limit of the detector, are approximately 0.5 order of magnitude stronger than those based on the astrophysical background model. In this case, our observational results surpass those of most previous studies, providing stronger constraints than all others in the mass range above $10^{17}$ g. The exclusion limit for PBHs as dark matter has also been extended to $4\times 10^{17}~\mathrm{g}$, which aligns with the best upper limit presented in Ref. Berteaud:2022tws. When comparing the ABM given by Eq. (11) with the PL model provided by Eq. (13), we found that they agree well within the range of observational data. However, when extrapolated beyond the observational range, the predictions of the PL model are significantly lower. This indicates that the PL model may suffer from overfitting, which could lead to an overly optimistic interpretation of the final constraints. Due to the lack of sufficient observational data, we are unable to employ a ${\chi}^{2}$ estimation method to account for the influence of background diffuse sources. Therefore, we believe that with adequate observational data, Insight-HXMT will be capable of providing even more stringent constraint results.
IV Conclusion
@@
In this work, we primarily utilize observational data from Insight-HXMT to improve the existing constraints on the abundance of PBHs in the mass range $2\times 10^{16}$$5\times 10^{17}~\mathrm{g}$. We first consider the photon flux emitted by PBHs within this mass range, incorporating contributions from both Galactic and extragalactic sources. In this range, the photon flux generated by positron annihilation due to Hawking radiation cannot be neglected, particularly when the black hole temperature falls below 511 keV, where annihilation photons dominate. Therefore, we account for both the photon flux directly produced by Hawking radiation and the photons resulting from positron annihilation.
@@
First, when we directly use the observational data from Insight-HXMT to derive the constraint results, due to the lack of high-energy observational data and the absence of background diffuse source subtraction, our results are weaker than those of previous studies. When we introduce the astrophysical background model, the constraints improve by approximately 1.5 orders of magnitude relative to the first scenario. Although the constraint results have improved significantly, we still do not have an advantage compared to previous studies.
@@
In our analysis, the energy range of the data used spans from 1 to 120 keV, whereas the actual upper energy limit of Insight-HXMT observations extends up to 250 keV. Although we used the PL model to estimate the constraints based on data from 1 to 250 keV, the results remain less precise. Moreover, due to the lack of actual observational data, we are unable to remove the impact of diffuse background sources. This limitation prevents us from obtaining tighter constraints when using the power-law model. We need more observational data, especially in the 120–250 keV energy range, to improve our results. Furthermore, we applied an existing astrophysical background model without conducting a detailed analysis or disentangling individual sources within the data. Consequently, the results obtained are not the most accurate. We anticipate that by addressing these issues in future work, we will be able to achieve more accurate and robust results.
Acknowledgements.
We thank Sai Wang, Yichao Li, and Chen Zhang for helpful discussions. We also thank Rui Huang for providing the observational data. This work was supported by the National SKA Program of China (Grants Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (Grants Nos. 12473001 and 12533001), the China Manned Space Program (Grant No. CMS-CSST-2025-A02), and the 111 Project (Grant No. B16009).